OpenStudy (anonymous):

| x -1| -2 =| 2x -5 | solve

OpenStudy (anonymous):

you have to either graph or work in cases

OpenStudy (anonymous):

if \(x>\frac{5}{2}\) then \(|x-1|=x-1,|2x-5|=2x-5\) solve \[x-1+2=2x-5\]

OpenStudy (anonymous):

if \(x<1\) then \(|x-1|=1-x,|2x-5|=5-2x\) solve \[1-x-2=5-2x\]

OpenStudy (anonymous):

and finally the third case is if \(1<x<\frac{5}{2}\) i will leave that one to you